# VideoChat-Flash: Hierarchical Compression for Long-Context Video Modeling

### v2

@@ Abstract
Long-context modeling is a critical capability for multimodal large language models (MLLMs), enabling them to process long-form contents with implicit memorization. Despite its advances, handling extremely long videos remains challenging due to the difficulty in maintaining crucial features over extended sequences. This paper introduces a Hierarchical visual token Compression (HiCo) method designed for high-fidelity representation and a practical context modeling system VideoChat-Flash tailored for multimodal long-sequence processing. HiCo capitalizes on the redundancy of visual information in long videos to compress long video context from the clip-level to the video-level, reducing the compute significantly while preserving essential details. VideoChat-Flash features a multi-stage short-to-long learning scheme, a rich dataset of real-world long videos named LongVid, and an upgraded “Needle-In-A-video-Haystack” (NIAH) for evaluating context capacities. In extensive experiments, VideoChat-Flash shows the leading performance on both mainstream long and short video benchmarks at the 2B and 7B model scale. It firstly gets 99.1% accuracy over 10,000 frames in NIAH among open-source models.
@@ 1 Introduction
Long-context video modeling is one of the most essential capabilities in multimodal large language models (MLLM). With it, MLLMs can handle hours of movies or documentaries, online long video streams, and so on using million-level video token processing.
@@
Through video-language alignment and instruction tuning, existing video MLLMs [58, 59, 29, 31, 72, 13, 78, 34, 65, 27, 4, 26] have made significant progress in minute-level video understanding, including video captioning and question answering. However, it remains highly challenging to build MLLMs for processing extremely long videos (lasting for hours or even longer). The most crucial difficulty lies in how to enable MLLMs to efficiently understand the multimodal long context brought by long videos.
@@
Inspired by large language models with long context, modeling multimodal long context is widely studied from several perspectives. Some work [47, 66] represented by Gemini-1.5-Pro [47] address it by training well-performed MLLMs on long-form corpus e.g. lengthy text and thousands of frames from videos, minimizing the gap between the evaluation and learning. To make the long training and inference work possibly and efficiently, building an effective training system with multimodal long context parallelism and data packing is crucial. Existing works evolve from the basic tensor parallelism to ring-style attention implementation and to current more comprehensive solutions integrating several advanced techniques. Although the progress in system construction and hardware has made it possible to train and infer with super-long multimodal contexts, such super-long multimodal contexts have significantly reduced the training and inference efficiency of models. (For Gemini-1.5-Pro [47], a one-hour video will be converted into 921,600 tokens). Meanwhile, the high redundancy in visuals makes it particularly difficult for models to understand. Some previous efforts [52, 15, 63] have been made to compress video tokens in order to achieve higher training and inference efficiency for long videos. Besides of these accomplishments, challenges still remain that increasing (or even maintaining) context understanding is often compromised by extending its length because compression on visuals trades off partial detail representations more or less, while the data and learning discrepancies between inference and training are hard to neutralize. Consequently, some current long video models even perform worse than some image-based MLLMs on video understanding benchmarks.
In this paper, we propose a Hierarchical video token Compression method (HiCo) with high-fidelity as well as a practical context modeling system VideoChat-Flash for multimodal long sequence processing. For HiCo, it widens context window using high-density token representations by compressing tokens hierarchically. We sequentially handle long videos, segmenting them into short clips, and feed clips into a MLLM by order. Similar to previous studies, HiCo condenses the visual tokens extracted from a clip to a few one based on spatiotemporal redundancies first. Linking these compressed visual tokens from clips and user query tokens together, HiCo then progressively reduce token quantity when going deeper in LLM processing, exploiting semantic correlations between clips and world knowledge embedding for compact representations.
With HiCo, we design a video MLLM VideoChat-Flash with a multi-stage short-to-long learning scheme and a corresponding data receipt with real-world long videos. In the training strategy, we start with short videos, and their captions and question-answers (QAs) for supervised fine-tuning, shift to long ones and their textual annotations, and finally employ a mixed corpus with both short and long videos for learning. We note short videos not only enhance basic visual perception, but also help MLLMs express extremely long visuals more concisely and precisely using compression, as validated in Sec 4.3. In data, we give a new long video dataset for supervised fine-tuning (SFT) in 300,000 hours and 2 billion words of textual annotations, in addition to previous ones [57, 66, 50]. Also, we extend the current “needle-in-a-haystack” (NIAH) from single-hop to multi-hop to ensure testing results correlate with models’ context modeling other than luck.
We contribute to the long video understanding as follows:
$\bullet$ We propose a new hierarchical compression scheme to encode long videos using much fewer tokens compared to existing methods. It addresses both visual and semantic redundancies, achieving a compute reduction by two orders of magnitude without compromising its long video understanding. In evaluations, our 7B model yields state-of-the-art performance with 99.1% retrieval accuracy over 10,000 frames in the typical “Needle-in-a-haystack” among open-sourced MLLMs. Also, it outperforms other 7B models in both short and long video evaluation benchmarks.
$\bullet$ We give an effective long video understanding MLLM VideoChat-Flash. It is trained with a three-stage short-to-long strategy. To support its scalable learning, we construct the corresponding long video data for supervised fine-tuning. Our training infrastructure includes a distributed system that supports high-degree sequence parallelism, ensuring efficient and robust model training.
$\bullet$ To concisely evaluate multimodal context capacities, we extend current NIAH in videos to a multi-hop configuration. It demands MLLMs to locate the target via a chain of visual hints and answer it, characterizing models’ context memorization and understanding rigorously.
@@ 2 Related Works
Figure 3: Framework of VideoChat-Flash with hierarchical token compression (HiCo).
Multimodal Large Language Models for Long Video Understanding.
Recent advancements in multimodal large language models (MLLMs) [47, 52, 63, 57, 38, 50, 66, 51] have shown significant promise in long video understanding. To address the challenge of processing long videos, researchers focus on two key strategies: (1) extending the context window of the LLM [47, 75, 66, 62] and (2) compressing the video tokens [33, 15, 63, 53, 52, 51]. For context extension, LongVILA [66] proposes a parallel processing system that distributes computation across temporal and tensor dimensions to optimize load balancing and inter-node communication. Although the approach of expanding the context window enable the possibility of long video understanding, it falls short of reducing the high computational burden and processing costs induced by long videos, thereby imposing limitations on its practical application. For token compression, Methods represented by Llama-Vid [33] use a highly compact representation while preserving key information. The high compression ratio makes it difficult for such methods to achieve excellent long video understanding performance, and they may even be inferior to some MLLMs designed for image modeling.
Long Video Benchmarking.
The evaluation of long video understanding remains a significant challenge, given the complex interplay of object recognition, temporal reasoning, memory, and higher-order cognitive abilities. Several benchmarks [46, 52, 73, 21, 71, 6, 19, 14, 32] and evaluation methodologies have emerged to address this need, primarily focusing on question-answering (QA) tasks. EgoSchema [40] leverages the Ego4D dataset to create a benchmark of 5000 multiple-choice QA questions on three-minute video clips. VideoMME [16] uses a collection of diverse online videos with human-annotated QA pairs. MLVU [81] expands the scope to include movies, TV dramas, and surveillance footage, evaluating performance on tasks such as temporal order reasoning, plot summarization, and QA. Focusing on fine-grained information retrieval, Video-NIAH [80] presents a “needle in a haystack” challenge, emphasizing the retrieval of specific details from long videos. Similarly, LongVideoBench [64] assesses both perceptual and reasoning abilities on fine-grained features within long video sequences.
@@ 3.1 HiCo: Efficient Long Video Modeling
To handle more than thousands of frame inputs, we approach the efficient long video understanding design from the perspective of multimodal context compression in MLLM, named as hierarchical compression (HiCo). We implement HiCo in a new MLLM VideoChat-Flash. As illustrated in Fig. 3, VideoChat-Flash leverages a three-component structure: a visual encoder, a visual-language connector, and a Large Language Model (LLM). The pipeline is divided into two main stages, specifically Clip-level Compression during the encoding of long videos and Video-level Compression within the context interaction in the LLM.
Dynamic Video Sampling.
Firstly, we need to perform frame sampling on the original video. We adopt a dual sampling strategy that both helps the model develop both micro and macro-level video understanding: fine-grained motion recognition in short clips, and broader event comprehension in longer sequences. For short videos, dense frame sampling (e.g. 15 fps) is taken to capture detailed motions. For long videos, sparse frame sampling (e.g. 1 fps) is taken to focus on event understanding.
Timestamp-aware Prompt
@@
For video MLLMs, the ability to perceive timestamps is also a crucial capability. Unlike previous works [48, 69] that rely on additional modules or designs to achieve this, we employ a simple timestamp-aware text prompt after the video context: “The video lasts for N seconds, and T frames are uniformly sampled from it.” We find that this straightforward approach is sufficient to enable the model to perceive the timestamps of the input video, achieving excellent performance on temporally sensitive tasks such as temporal grounding. (see Tab. 3 for details)
Encoding and Compression for Video Clips.
@@
Considering the abundant redundant and repetitive information such as backgrounds and objects that exists between adjacent frames in natural videos, we sequentially encoding a long video clip by clip, and compress its visual representation through inter-frame redundancies and semantic correlations. Formally, given a frame sequence sampled from the original video, we segment it into $T$ equally sized clips. The frames of $j_{\text{th}}$ clip are transformed by a visual encoder $\mathcal{V}$ into $M$ visual tokens, denoted as $[\mathbf{v}_{i}^{j}]_{i=1,2,..,M}$. We design a video-language connector $\mathcal{F}$ composed of a compressor and an MLP projection. It compresses $M$ tokens into $N$ ones, $[\mathbf{s}_{i}^{j}]_{i=1,2,..,N}$ (w.r.t. $N<M$), before input to a LLM:
$[\mathbf{s}_{i}^{j}]_{i=1,2,..,N}=\mathcal{F}([\mathbf{v}_{i}^{j}]_{i=1,2,..,M}),$ (1)
where $\mathcal{F}$ denotes our given connector. The specific design of the visual encoder and compressor is a non-trivial decision as lots of effective candidates (e.g. Video with Q-former, Image encoders with MLP, etc.) have been explored [31, 78]. We systematically study how visual encoders and connectors are combined in different cases for the best spatiotemporal compressed representation. As shown in Fig. 4, we consider three different numbers of tokens per frame after compression (16, 49, 196) and examined image encoders, video encoders, and four popular token compression strategies: spatial downsampling [78, 12], uneven downsampling [62], spatio-temporal resampler [31, 53], and similar token merging [5, 63]. We found that video encoder, by modeling inter-frame interactions, better exploit temporal redundancy compared to image encoder. Consequently, combining a video encoder with a parameter-free similar token merge operation facilitates learning effective spatiotemporal compressed representations.
Figure 4: Comparison of different compressors. The specific comparisons and implementation details can be found in the Sec. 4.3 and the Appendix.
@@
Although clip-level compression has been carried out before, due to the possibility of longer-range visual redundancies in long videos (e.g. surveillance videos), and when an LLM responds to specific instructions regarding the visual input, it may not be necessary to continuously focus on the entire long video context. We consider conducting further video-level compression during the LLM inference stage. Recent works [9, 8] have explored acceleration strategies for MLLMs when processing short visual contexts. Most of them drop visual tokens based on the correlation between text tokens and visual tokens. In contrast, we have found that when an LLM processes a long video context, it pays attention to the entire long video context at the shallow layers of the LLM, while focusing on the details of certain local moments at the deep layers (see the Appendix for specific visualizations). Based on this observation, we have designed a progressive visual dropout strategy, which is divided into two stages. At the shallow layers of the LLM, we uniformly drop a small number of video tokens (i.e. uniform drop), reducing the computation while maintaining the original spatio-temporal structure of the video context. At the deep layers of the LLM, we rely on the correlation between text tokens and video tokens to retain the most critical relevant information (i.e. text-guided select). With the $NT$ visual tokens $[\mathbf{s}_{i}]^{j}$ from the long video, we further condense them in a progressive manner during the computation in LLM by dropping visual ones. Formally, we denote $[\mathbf{t}_{i}^{(0)}]_{i=1,2,..,NT}=[\mathbf{s}_{1}^{1},\mathbf{s}_{2}^{1},...,\mathbf{s}_{1}^{T},\mathbf{s}_{2}^{T},...,\mathbf{s}_{N}^{T}]$ as the input visual tokens of the LLM, and $[\mathbf{t}_{i}^{(l)}]$ as the $l_{\text{th}}$ output of these visual tokens via LLM. Regarding this,
$[\mathbf{t}_{i}^{(l)}]=[\mathbf{r}_{i}^{(l)}*\mathbf{t}_{i}^{(l-1)}]\quad\text{w.r.t.}\quad\mathbf{r}_{i}^{(l)}\in\{0,1\}$ (2)
We have found that this operation not only effectively improves the computational efficiency of the model but also enhances the understanding performance of the model by reducing irrelevant visual noise.
@@ 3.2 Multi-stage Short-to-Long Learning
We focus on using videos and their annotations for training. Unlike previous works [75, 66] that often use long-form text to extend the context window of LLMs, we argue that direct training on long-form videos minimizes the gap between training and testing, leading to better performance in downstream long video evaluations. Using a short-to-long scheme, our proposed VideoChat-Flash is trained on a mixed dataset of both short and long videos. The training data are detailed in the Appendix.
@@
To enable the model to handle a wide variety of video tasks, we collect 3.5 million instruction fine-tuning samples, including 1.1M images, 1.7M short videos (under 60 seconds), and 0.7M long videos (60$\sim$3600 seconds). We mix the short and long video data to ensure the model retains fine-grained understanding while expanding its comprehension of long videos. The sampling method used is the dynamic video sampling described in Section 3.1, with the number of sampled video frames ranging from 64 to 512.
@@
3.3 Large-scale Corpus for Long Video Training
Our proposed dataset leverages the rich diversity of existing datasets, including Ego4D [18], HowTo100M [41], HD-Vila [67], and MiraData [23], encompassing a wide range of video types: movies, egocentric videos, news, interviews, and how-to videos, and other in-the-wild videos of long duration. For data curation, we generate dense event labels for each long video. Specifically, we utilize existing high-quality short video captions (Panda-70M [10] for HD-VILA [67], CosMo [54] for HowTo100M [41], Ego4D-HCap [22] for Ego4D [18], and the original high-quality captions provided in MiraData [23]) and filter the consecutive segments that can be regroup into a long video sequence, then we construct a sequence of event labels with their corresponding timestamps for every long video based on their captions. In this process, for datasets with high-quality event-level annotations (HT-Step [1] for HowTo100M [41], Ego4D-HCap [22] for Ego4D [18]), we directly utilize them as the event labels, while for others, we extract the major event from the caption using an LLM. Finally, we construct several types of long video QA pairs based on the video captions, event labels, and the timestamps of short video segments.
The annotations of these videos are categorized into five tasks: video captioning, temporal grounding, event relation recognition, scene relation recognition, and video event counting. Among them, the temporal grounding task is to answer the accurate timestamp for one specific event, and event relation recognition is constructed by selecting the right event order in a long video. Scene relation recognition is formulated analogously to the event relationship one. It is built for videos that have background captions (like MiraData), where we can extract the major scenery labels.

### v4

@@ Abstract
Long-context video modeling is critical for multimodal large language models (MLLMs), enabling them to process movies, online video streams, and so on. Despite its advances, handling long videos remains challenging due to the difficulty in efficiently understanding the extremely long video context. This paper aims to address this issue from aspects of the model architecture, training data, training strategy and evaluation benchmark. First, we propose a novel Hierarchical video token Compression (HiCo) method, which leverages visual redundancy in long videos to compress long video context from Clip-level to Video-level, reducing the computation significantly while preserving essential details, achieving an extreme compression ratio of approximately 1/50 with almost no performance loss. Second, we introduce a multi-stage short-to-long learning scheme, a large-scale dataset of real-world long videos named LongVid, and a challenging “Multi-Hop Needle-In-A-Video-Haystack” benchmark. Finally, we build a powerful video MLLM named VideoChat-Flash, which shows a leading performance on both mainstream long and short video benchmarks at the 2B and 7B model scale. It first gets 99.1% accuracy over 10,000 frames in NIAH among open-source models.
@@ 1 Introduction
Long-context video modeling stands as one of the most crucial capabilities within multimodal large language models (MLLMs). This capability empowers MLLMs to proficiently manage hours-long movies, documentaries, and online video streams, all of which demand sophisticated long video processing.
@@
Recent advances in MLLMs [55, 56, 27, 29, 66, 14, 70, 31, 60, 25, 5, 24] perform well in short video understanding. However, it remains challenging to build MLLMs for processing extremely long videos (lasting for hours or even longer). The difficulty lies in how to enable MLLMs to efficiently understand the extremely long video context brought by long videos.
@@
Inspired by large language models (LLMs) with long context, modeling multimodal long context is widely studied from several perspectives. Some work [44, 61] represented by Gemini-1.5-Pro [44] address it by training well-performed MLLMs on long-form corpus e.g. lengthy text and thousands of frames from videos, minimizing the gap between the evaluation and learning. Although the progress in system construction and hardware has made it possible to train and infer with super-long multimodal contexts, such super-long multimodal contexts have significantly reduced the training and inference efficiency of models. (For Gemini-1.5-Pro [44], a one-hour video will be converted into 921,600 tokens). Meanwhile, the high redundancy in long video context makes it particularly difficult for models to understand. Some previous efforts [49, 30, 47] have been made to compress video tokens in order to achieve higher training and inference efficiency for long videos. However, the compression of visual content inevitably leads to the loss of detailed information. In some long video understanding benchmarks, certain current long video models even perform worse than some image-based MLLMs. Therefore, how to strike a balance between performance and efficiency remains a significant challenge. In this paper, we attempts to address the above issues from the model architecture, training data, training strategy and evaluation benchmark.
First, we propose a novel Hierarchical video token Compression method (HiCo) to model the long video context efficiently, which defines the compression of the long video context into two stages. First, we segment the long video into multiple clips. Then, at the Clip-level, we utilize the spatio-temporal attention of the video encoder and the similar token merging to aggregate the key information between frames, thereby reducing the redundancy of inter-frame features. Subsequently, we take advantage of the sparsity of attention when the LLM processes long video tokens, further discard the video tokens that are irrelevant to the current task at the Video-level. HiCo could achieve an extreme compression ratio of approximately 1/50 with almost no performance loss. Additionally, we have conducted thorough explorations of other designs such as video sampling and timestamp awareness prompt.
Second, to further enrich the existing long video training corpus, we construct LongVid, a dataset that contains 300,000 hours of videos and 2 billion words of textual annotations. With LongVid, we have designed a multi-stage training strategy named short-to-long learning. The main idea is to first utilize image and short video data to learn basic visual perception abilities. Then, through the joint training of short video and long video data, the model is enabled to handle videos of different lengths and different types of tasks. In addition, we design a new evaluation benchmark named “Multi-Hop Needle In A Video Haystack”. Which is more challenging and can better examine the model’s complex reasoning abilities regarding long videos.
Finally, we develop a powerful video MLLM named VideoChat-Flash, as shown in Fig. 2, which achieves remarkably leading performance with extremely high efficiency on various video understanding benchmarks. Even with a 7B size, it outperforms closed-source models such as GPT-4o [40] and Gemini-1.5-Pro [44]. And it first yields 99.1% retrieval accuracy over 10,000 frames in the “Needle-In-a-Video-Haystack” among open-sourced MLLMs.
@@ 2 Related Works
Figure 3: Framework of VideoChat-Flash with Hierarchical Video Token Compression. Video tokens will be compressed at the Clip-level by leveraging the local redundancy of the video modality during video encoding. Subsequently, during LLM processing, they will be compressed at the Video-level by taking advantage of the sparsity in the interaction between the text modality and the video modality.
Multimodal Large Language Models for Video Understanding.
Recent advancements in multimodal large language models (MLLMs) have shown significant promise in video understanding. Most of them [27, 29, 56, 31, 66, 70, 71] focus on the understanding of minute-level videos, and some works [44, 49, 54, 47, 61, 48, 20] have further tried to handle longer hour-level videos. To address the challenge of processing long videos, researchers focus on two key strategies: (1) extending the context window of the LLM [44, 68, 61, 57] and (2) compressing the video tokens [30, 15, 58, 50, 49, 48, 64]. For context extension, although the approach of expanding the context window enable the possibility of long video understanding, it falls short of reducing the high computational burden and processing costs induced by long videos, thereby imposing limitations on its practical application. For token compression, Methods represented by Llama-Vid [30] use a highly compact representation while preserving key information. The high compression ratio makes it difficult for such methods to achieve excellent long video understanding performance, and they may even be inferior to some MLLMs designed for image modeling. Therefore, how to design a Video MLLMs architecture that can balance both efficiency and performance remains a difficult challenge. In this work, we provide a comprehensive solution that balances both efficiency and performance from various aspects such as the model architecture, training data, and training strategies.
Long Video Benchmark.
In order to evaluate the ability of Video MLLMs to understand long videos, previous works [49, 67, 36, 43, 59, 73, 16, 7, 53] have achieved this by collecting long videos and then designing various multiple-choice questions related to the content of these long videos. This approach is closer to real-world applications and can effectively examine the model’s ability to understand and reason about long videos. However, when it comes to examining the model’s capabilities for videos of different lengths, this method is not intuitive enough. Inspired by the popular ”Needle in A Haystack” (NIAH) evaluation in long text context evaluation, some recent works [68, 72] have attempted NIAH for Video haystack. Nevertheless, it is difficult to assess complex reasoning abilities, and there may be information leakage. In this paper, we propose a more challenging “Multi-Hop Needle-In-A-Video-Haystack” is designed to address the above issues.
@@ 3.1 HiCo: Efficient Long Video Modeling
To enable MLLMs to handle thousands of input frames, we propose a new video context compression paradigm named hierarchical compression (HiCo). This paradigm decomposes video context compression into two main stages: 1. Clip-level Compression during the encoding of long videos. 2. Video-level Compression within the context interaction in the LLM. Based on this framework, we have designed an innovative efficient Video MLLM architecture, VideoChat-Flash, as illustrated in Fig. 3. Below, we elaborate on our specific design details from data input to model output.
Duration-based Sampling.
First, we need to perform frame sampling on the original video. Specifically, we sample a raw video with a duration of $D$ to obtain $T$ frames as input. Considering that the requirements for understanding short and long videos often differ, we aim to conduct dense sampling on short videos to capture detailed motions and sparse sampling on long videos to focus on event understanding. To this end, we have designed a Duration-based Sampling strategy:
$T=\min(T_{\text{max}},\max{(D,T_{\text{min}})}).$ (1)
Simultaneously, we define the sampling density $\phi$ as follows:
$\phi(T,D)=\frac{T}{D}=\frac{\min(T_{\text{max}},\max{(D,T_{\text{min}})})}{D}.$ (2)
That is, for short videos where $D<T_{\text{min}}$, $\phi=T_{\text{min}}/D$ , which increases as the video length decreases. For long videos where $D>T_{\text{max}}$, $\phi=T_{\text{max}}/D$, which decreases as the video length increases.
Timestamp Prompt.
@@
For video MLLMs, the ability to perceive timestamps is also a crucial capability. Unlike previous works [45, 64] that rely on additional modules or designs to achieve this (there is a considerable computational burden when there are a large number of video frames), we employ a simple timestamp prompt after the video context: “The video lasts for N seconds, and T frames are uniformly sampled from it.” We find that this straightforward approach is sufficient to enable the model to perceive the timestamps of the input video, achieving excellent performance on timestamp sensitive tasks such as temporal grounding (see Tab. 1).
Spatio-Temporal Compression Encoding for Clips.
@@
Considering the substantial redundant and repetitive information, such as that of backgrounds and objects, present between adjacent frames in natural videos, we segment the original video frames into several clips. Subsequently, we employ a video encoder with spatio-temporal attention to encode these clips. This enables each visual token to aggregate information from other frame tokens as comprehensively as possible. Finally, we utilize token merging to combine highly similar tokens. Formally, given a frame sequence sampled from the original video, we divide it into $N_{c}$ equally sized clips. The frames of $j_{\text{th}}$ clip $\mathbf{x}^{j}$ are transformed by a video encoder and a connector $\mathcal{F}$, resulting in $M$ compressed visual tokens:
$[\mathbf{v}_{i}^{j}]_{i=1,2,..,M}=\mathcal{F}(\mathcal{V}(\mathbf{x}^{j})),$ (3)
where $\mathcal{F}$ consists of a parameter-free similar token merge operation and an MLP projection. Ultimately, we concatenate the compressed tokens of each clip to obtain the input for the LLM:
$\mathbf{X_{v}}=\text{Concat}([\mathbf{v}_{i}^{1}]_{i=1,2,..,M},\cdot\cdot\cdot,[\mathbf{v}_{i}^{N_{c}}]_{i=1,2,..,M}).$ (4)
Benefiting from the effectiveness of the video encoder in modeling spatio-temporal interactions, we achieve an extremely heavy compression while well retaining the key information, with each video frame being compressed to an average of only 16 tokens.
@@
Although clip-level compression has been carried out before, due to the possibility of longer-range visual redundancies in long videos (e.g. surveillance videos), and when an LLM responds to specific instructions regarding the visual input, it may not be necessary to continuously focus on the entire long video context. We consider conducting further video-level compression during the LLM inference stage. Recent works [10, 9] have explored acceleration strategies for MLLMs when processing short visual contexts. Most of them drop visual tokens based on the correlation between text tokens and visual tokens. In contrast, we find that when the LLM processes a long video context, it pays attention to the entire long video context at the shallow layers of the LLM, while focusing on the details of certain local moments at the deep layers (see the Appendix for specific visualizations). Based on this observation, we have designed a progressive visual dropout strategy, which is divided into two stages. At the shallow layers of the LLM, we uniformly drop a small number of video tokens (i.e. uniform drop), reducing the computation while maintaining the original spatio-temporal structure of the video context. At the deep layers of the LLM, we rely on the correlation between text tokens and video tokens to retain the most critical relevant information (i.e. text-guided select). We have found that this operation not only effectively improves the computational efficiency of the model but also slightly enhances the understanding performance of the model by reducing irrelevant visual noise.
3.2 Large-scale Corpus for Long Video Training
To build LongVid, we leverages the rich diversity of existing datasets, including Ego4D [19], HowTo100M [37], HD-Vila [62], and MiraData [22], encompassing a wide range of video types: movies, egocentric videos, news, interviews, and how-to videos, and other in-the-wild videos of long duration. For data curation, we generate dense event labels for each long video. Specifically, we utilize existing high-quality short video captions (Panda-70M [11] for HD-VILA [62], CosMo [51] for HowTo100M [37], Ego4D-HCap [21] for Ego4D [19], and the original high-quality captions provided in MiraData [22]) and filter the consecutive segments that can be regroup into a long video sequence, then we construct a sequence of event labels with their corresponding timestamps for every long video based on their captions. In this process, for datasets with high-quality event-level annotations (HT-Step [1] for HowTo100M [37], Ego4D-HCap [21] for Ego4D [19]), we directly utilize them as the event labels, while for others, we extract the major event from the caption using an LLM. Finally, we construct several types of long video QA pairs based on the video captions, event labels, and the timestamps of short video segments.They are categorized into five tasks: video captioning, temporal grounding, event relation recognition, scene relation recognition, and video event counting. See Appendix for more details.
@@ 3.3 Multi-stage Short-to-Long Learning
Unlike studies [68, 61] that use long-form text to extend the context window, we prefer that direct training on long-form videos minimizes the gap between training and testing, leading to better downstream evaluations. Using a short-to-long scheme, our proposed VideoChat-Flash is trained on a mixed dataset of both short and long videos. The training data are detailed in the Appendix.
@@
To enable the model to handle a wide variety of video tasks, we collect 3.5 million instruction fine-tuning samples, including 1.1M images, 1.7M short videos (under 60 seconds), and 0.7M long videos (60$\sim$3600 seconds). We mix the short and long video data to ensure the model retains fine-grained understanding while expanding its comprehension of long videos. The sampling method used is the duration-based sampling described in Section 3.1, with the number of sampled video frames ranging from 64 to 512.